Problem 5:20 points) Find the force P to move Block Aupwards if block A has a mass of 40 kg and block B has a mass of 35 kg. For friction, .25 and 2everywhere. 15 15 Need help please
Answer #1

Given: mA= 40 kg , mB= 35 kg.

\muk= 0.25 and \mus= 0.2

Block A will move upwards only when block B will move forward.

Hence modeling this problem as forward moving of block B.

Friction force acting on block B , that will restrict to move :

Fs=\mus [(mB.g sin15) + (mA.gsin15)]

= 0.2[ (35×9.81×sin15) + (40×9.81×sin15)]

= 0.2[88.87 + 101.56]

= 38.086 N

Fk= 0.25[88.87+101.56]

= 47.61 N

Force required to move , P = Fs+Fk = 38.086+47.61 = 85.7 N